Fermi National Laboratory
 

Operation Summary for August 5, 2011 - August 8, 2011

 

The Integrated Luminosity for the period from 8/1/11 to 8/8/11 was 5.82 pb-1 (inverse picobarns). During the same time period NuMI took 6.57E18 of beam.

Friday August 5

The day shift began with the Tevatron (TEV) off for cryo systems and vacuum systems recovery, with the Booster, Main Injector (MI), and Recycler off and in access, with Operations monitoring the Antiproton Source (Pbar) stack (97.9E10), with the Linac operating normally and with the Neutron Therapy Facility (NTF) preparing to treat a patient, with the NuMI, MiniBooNE, and Meson FTBFT-1010 experiments off due to other accesses, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments off, in access, and conducting maintenance work.

During the day shift, MTAbegan taking beam. FESS personnel reported at 9:23 AM that they had completed maintenance on the MI LCW pumps at MI-30. Operators reported at 10 AM that an individual in MI-40 opened an emergency door and dropped the interlocks; operators searched and secured the enclosure. The Transfer Hall came out of access at 10:08 AM. NTF completed treating its only patient for the day at 10:46 AM. MI came out of access at noon. Booster came out of access at 12:15 PM. Operations sent beam to Pbar for stacking and to the NuMI and MiniBooNE experiments at 12:44 PM. MI personnel reported problems with beam positions in MI-8 and were also investigating damper problems at MI-60. Operators reported at 2:30 PM that they were preparing the TEV for access so that Vacuum personnel could fix a problem with a vacuum valve in TEV sector A4 (BV9D). The Cryo coordinator reported at 3:35 PM that the TEV cryo system was cold. Operations disabled beam to MTA at 4 PM.

During the evening shift, operators reported at 5:05 PM that there was a problem with the Tevatron Electron Lens (TEL1); operators notified TEL personnel. At 7:37 PM, operators prepared the TEV for access to deal with vacuum problems. Operations disabled the beam to T-1010 at 7:38 PM. Operators and a Vacuum person walked through TEV sector A1, opened vacuum valves, and came out of access at 8:41 PM. At 10:08 PM, during the TEV turn on, sector B2 quenched during the first ramp. The B2 cryo system had recovered by 11:16 PM and Operations had put the TEV through the first of three ramps at 11:59 PM.

Saturday August 6

The midnight shift began with the TEV ramping, with Operations monitoring the Recycler stash (159.3E10), with Pbar stacking (130.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operators put the TEV through a dry squeeze at 12:14 AM. Operators put the TEV through a wet squeeze at 2:11 AM. Operators called in a TEV Controls person at 2:19 AM to help with an E3 BPM communication problem. At 3:57 AM, a number of vacuum pumps for the Linac I- Source tripped off; they all reset. A TEV Controls person reported at 3:59 AM that he needed to replace the processor board for a timing signal generator. Operators began preparing the Pelletron for an access.

During the day shift, Operations established beam to T-1010. Operators reported at 10:19 AM, that Controls personnel had changed out the TEV sector E3 BPM chassis. TEV personnel conducted orbit checks and then put the TEV through a wet squeeze at 11:11 AM. Operations put the TEV into shot setup at 1:19 PM and established store 8952 at 2:57 PM with an average initial luminosity (AIL) of 74.93E30.

During the evening shift, Pelletron personnel reported at 5:35 PM that they found a damaged water pump pulley; no estimate for repairs. Operations disabled beam to T-1010 at 6:02 PM. At 9:18 PM, the MI Power Supply interlock Loop (MI PSLOOP) tripped off; it reset. Operations terminated store 8952 at 11:13 PM and put the TEV into shot setup.

Sunday August 7

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 8953 at 12:37 AM with an AIL of 129.1E30. At 3:49 AM, a Booster injection pulsed magnet (MP02) tripped off when its vacuum pump (IPMP02) tripped off for the first of two times in six minutes; operators investigated and reported that the vacuum pump’s power supply had failed. Operators called in a Booster person who had the supply replaced by 4:53 AM. At 5:08 AM, the P1 transfer line beam permit tripped off for the first of many times in one hour on a power supply high limit (I:HT701). Operators contacted the Operations Pbar Specialist who instructed them to expand the limits on HT701 and HT702. Operations established beam to T-1010 at 7:21 AM. A Booster RF station (BRF15) tripped off for the first of three times in three minutes; operators bypassed the station after the last trip and contacted a Booster person.

During the day shift, Operations terminated store 8953 at 8:37 AM, put the TEV into shot setup, and established store 8954 at 9:18 AM with an AIL of 128.1E30. At 11:25 PM, store 8954 aborted; there was no quench. Operators investigated and found a TEV anode power supply control rack at the F0 service building with out power; operators called in RF personnel. RF personnel traced the problem to the secondary side of the yard breaker; FESS personnel checked the breaker, could not find a problem, and reset it. Operators put the TEV through a wet squeeze. Operations put the TEV into shot setup at 3:02 PM.

During the evening shift, Operations established store 8957 at 4:09 PM with an AIL of 112.76E30. Pelletron personnel reported at 5:51 PM that they thought the repairs would be completed in about an hour and would want the SF6 gas pumped back into the Pelletron. Operations disabled beam to T-1010 at 6:01 PM. Operations terminated store 8957 at 11:08 PM and then prepared CDF for access.

Monday August 8

The midnight shift began with CDF in access, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, CDF came out of access at 12:15 AM and operators put the TEV through a dry squeeze. Operations put the TEV into shot setup at 1:33 AM, but the TEV aborted at 2:10 AM due to bad beam positions at TEV sector D0; operators and a TEV person investigated and reported that the E48 kicker prefired. Operations reported at 2:56 AM that there were still antiprotons left in the TEV that they could not abort. A TEV person came in, contacted D0, and had them make up their abort, which allowed Operations to abort the remaining beam. Operations established beam to T-1010 at 4:40 AM. Operations put the TEV into shot setup at 5:13 AM and established store 8961 at 6:13 AM with an AIL of 79.1E30. Operators reported that the transfer of SF6 gas into the Pelletron began at 7:33 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to work on their damper problems. MTA will take beam today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/08/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory